OpenStudy (anonymous):

Find derivative of (2x+1)^3 sin(4x)

OpenStudy (anonymous):

product and chain for this one

OpenStudy (anonymous):

\[\left(fg\right)'=f'g+g'f\]with \[f(x)=)2x+1)^3,f'(x)=3(2x+1)^2\times 2, g(x)=\sin(4x), g'(x)=4\cos(4x)\]

OpenStudy (anonymous):

some kind of typo there, but maybe it is clear

OpenStudy (anonymous):

\[f(x)=(2x+1)^3,f'(x)=3(2x+1)^2\times 2\]\[ g(x)=\sin(4x), g'(x)=4\cos(4x)\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

Why is there a x2 at the end?

OpenStudy (anonymous):

because the derivative of \(2x+1\) is \(2\)

OpenStudy (anonymous):

really it is \[f'(x)=6(2x+1)^2\]

OpenStudy (anonymous):

Ohh OK! Do you think you can help me with one more?

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

you still have to put all this together using the product rule, but that is just a matter of writing it

OpenStudy (anonymous):

you can post the next one here if you like

OpenStudy (anonymous):

OK. The other one is (2x+1)^3 tan(x)

OpenStudy (anonymous):

really? it is almost identical to this one again you need the product rule, which i hope you know \[\left(fg\right)'=f'g+g'f\]

OpenStudy (anonymous):

But doesn't the tan make it different?

OpenStudy (anonymous):

this time \[f(x)=(2x+1)^3,f'(x)=6(2x+1)^2\] \[g(x)=\tan(x), g'(x)=\sec^2(x)\] yeah it makes a difference, the derivative of tangent is secant squared, not cosine!

OpenStudy (anonymous):

but the formula and the work is still the same

OpenStudy (anonymous):

Got it from there. Thanks for your patience! Haha I'M kinda slow

OpenStudy (anonymous):

you do understand you have to put all that together in this \[\left(fg\right)'=f'g+g'f\] right?

OpenStudy (anonymous):

Less. My Cow assignment will only take it in that form

OpenStudy (anonymous):

slow schmo you were not born knowing this it all takes some practice then you can forget all about it later

OpenStudy (anonymous):

*yes

OpenStudy (anonymous):

COW as in temple's cow?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you go there? or just use the system?

OpenStudy (anonymous):

Our teacher has us just use the system

OpenStudy (anonymous):

oh wow it is ancient surprised they don't use mylabs plus or webassign hope dan reich is making some money out of this

OpenStudy (anonymous):

Well I do live in south Dakota aka the middle of nowhere so, figures. Haha